OpenStudy (anonymous):

@loser66 If sin y= cos 3y, determine an exact value for angle y.

OpenStudy (anonymous):

I think I just got one of many of the solutions for this question by fluke.

OpenStudy (anonymous):

Heres how I went about doing it. \[\sin y= \cos 3y\] \[\frac{ \sin y }{ \cos 3y }=1\] \[\tan (\frac{ 1 }{ 3 }y)=1\] \[\frac{ y }{ 3 }=\frac{ \Pi }{ 4 }\] \[y=\frac{ 3\Pi }{ 4 }\] Is my method correct or was it just a fluke?

OpenStudy (loser66):

nope, it is not valid , since \(\dfrac{siny}{cos 3y}\neq tan(y/3)\)

OpenStudy (anonymous):

yea I thought so Thanks for confirming

OpenStudy (anonymous):

so how would I go about doing this?

OpenStudy (loser66):

I am sorry, I don't have time for now.

OpenStudy (loser66):

@iambatman